a+b+c=7 1/(a+b) + 1/(b+c) + 1/(c+a)= 7/10 a/(b+a) + b/(c+a) + c/(a+b)=?

a+b+c=7
1/(a+b) + 1/(b+c) + 1/(c+a)= 7/10
a/(b+a)  + b/(c+a) + c/(a+b)=?

Grade:

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
please find the answer to your question below
Given a+b+c=7
so a=7-(b+c) ,b=7-(c+a) ,c=7-(a+b)..............(1)
also given 1/(a+b)+1/(b+c)+1/(c+a)=7/10.............(2)
so (a/(b+c))+(b/(c+a))+(c/(a+b)) is
=(7-(b+c)/(b+c))+(7-(c+a)/(c+a))+(7-(a+b)/(a+b)) from 1
=(7/(b+c))+(7/(c+a))+(7/(a+b))-3
=7(7/10)-3 from 2
=4.9-3
=1.9